OpenStudy (anonymous):

Differentiate Phi(t) = exp{ λ(e^t − 1)},

OpenStudy (anonymous):

with respect to t?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

does exp mean e^()

OpenStudy (anonymous):

yeps!

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

exp{ λ(e^t − 1)}, y = lambda. just easier to type = e^(y*e^t - y) break up the e = [e^(y*e^t)]*(e^-y)

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

you can type it into wolfram and see what it spits out. maybe that will help you work backward and see how they arrive at the answer.

OpenStudy (anonymous):

\[\large y=e^{y(e^t - 1)}\] \[\large \frac{dy}{dt}=e^{y(e^t - 1)}[(e^t-1)(\frac{dy}{dt})+y(e^t)]\] \[\large \frac{dy}{dt}=e^{y(e^t - 1)}(e^t-1)(\frac{dy}{dt})+e^{y(e^t - 1)}ye^t\] \[\large \frac{dy}{dt}-e^{y(e^t - 1)}\frac{dy}{dt}=e^{y(e^t - 1)}ye^t\] \[\large \frac{dy}{dt}[1-e^{y(e^t - 1)}]=e^{y(e^t - 1)}ye^t\] \[\large \frac{dy}{dt}=\frac{e^{y(e^t - 1)}ye^t}{1-e^{y(e^t - 1)}}\] \[\large \frac{dy}{dt}=\frac{ye^{y(e^t-1)+t}}{1-e^{y(e^t - 1)}}\]